C-C BOND ACTIVATION AND COUPLING OF PROPENE INDUCED BY LA ATOM

\underline{DILRUKSHI HEWAGE}, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055; HONG TAO, Department of Chemistry, Southwest Forestry University, Kunming 650224, PR China; RUCHIRA SILVA, SUDESH KUMARI, AND DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

A series of La(C_{n}H_{m}) complexes with n ≤ 6 and m ≤ 12 were produced by the reactions between propene and La in a supersonic molecular beam source. Their formation and structures were investigated using mass-analyzed threshold ionization (MATI) spectroscopy in combination with theoretical calculations. Previously, we identified the formation of La(C_{3}H_{4}) and H-La(C_{3}H_{5}) through dehydrogenation and metal insertion mechanisms. In this work, we will discuss the formation of La(CH_{2}) and La(C_{4}H_{6}) by La induced C-C bond activation and coupling. La(CH_{2}) is formed by the C-C bond breakage and 1,2-hydride shift of propene and is a Schrock-type carbene complex. This complex is then coupled with the C=C bond of a second propene molecule to form La(C_{4}H_{6}) by removing two hydrogen atoms. The resultant La(C_{4}H_{6}) complex was idetified in two low-energy isomeric forms: one was a metallacycle (isomer A) and the other was lanthanum trimethylenemethane (isomer B). Both La(C_{4}H_{6}) isomers are in a doublet ground state, with isomer A in C_{s} point group and isomer B in C_{3}_{v}. Adiabatic ionization energies and several vibrational frequencies of the two complexes were obtained from the sharp MATI spectra.